OpenStudy (arbyj):

if the slope is -1 with points (-5, -11), then what's the equation? I've worked it out but I keep getting it wrong, help?

OpenStudy (shamil98):

Input your values of m, and (x,y) into \[\huge y-y_1=m(x-x_1)\]

OpenStudy (arbyj):

i forgot to add it has to be in slope intercept form

OpenStudy (shamil98):

Yes, input your values, and solve for y.

OpenStudy (shamil98):

\[\huge y + 11= -1(x+5)\] \[\huge y + 11 = -x - 5\] \[\huge y = -x - 16\]

OpenStudy (arbyj):

Oh okay I see, thank you

OpenStudy (shamil98):

np